planet X travels in a circular motion orbit around the Sun. The radius of Planet X is twice that of Earth. The year on Planet X is 540 Earth years. The mass of Planet X is half that of Earth. The mass of the sun is 1.99*10^30 kg. The Earth's orbital distance from the Sun is 1.50*10^11 m.
Answer #1

what we have to find is not clearly mentioned in the question thats why as per my knowledge i had finded out the orbital distance of that arbitrary planet X from the sun here R1=the orbital distance of earth from sun and R2= the orbital distance of planet X fom sun

the solution has bben done usingthe formula of the time period

so after calculating we get R2/R1=1/8.4 which we can take as 1/8 so R2=(1/8)R1